# Black hole X-ray binary A062000 in quiescence: hints of Faraday rotation of near-infrared and optical polarization?

### v1

@@
Black hole X-ray binary A062000 in quiescence: hints of Faraday rotation of near-infrared and optical polarization?
We present simultaneous high-precision optical polarimetric and near-infrared (NIR) to ultraviolet (UV) photometric observations of low-mass black hole X-ray binary A062000 in the quiescent state. Subtracting interstellar polarization, estimated from a sample of field stars, we derive the intrinsic polarization of A062000. We show that the intrinsic polarization degree (PD) is variable with the orbital period with the amplitude of at least in the band, where the signal-to-noise ratio of our observations is the best. It implies that some fraction of the optical polarization is produced by scattering of stellar radiation off the matter that follows the black hole in its orbital motion. In addition, we see a rotation of the orbit-average intrinsic polarization angle (PA) with the wavelength from in the to in the band. All of the above, combined with the historical NIR to optical polarimetric observations, shows the complex behavior of average intrinsic polarization of A062000 with the PA making continuous rotation from infrared to blue band by in total, while the PD remains nearly constant over the entire spectral range. The spectral dependence of the PA can be described by Faraday rotation with the rotation measure of RM= rad m-2, implying a few Gauss magnetic field in the plasma surrounding the black hole accretion disk. However, our preferred interpretation for the peculiar wavelength dependence is the interplay between two polarized components with different PAs. Polarimetric measurements in the UV range can help distinguishing between these scenarios.
@@
A062000 is the prototypical low-mass X-ray binary (LMXB) system containing a black hole (BH), discovered during its 1975 outburst (Elvis et al., 1975). Its luminosity increased by more than a million times compared to the quiescent levels reaching the Eddington luminosity. Since then, the object has resided in the quiescent state, where its spectrum is dominated by the emission of K-type star orbiting a BH (McClintock & Remillard, 1986; van Grunsven et al., 2017) with the period of h. Nevertheless, an additional source of non-stellar origin is needed to describe the excess of radio, near-infrared (NIR), and ultraviolet (UV) emission of A062000 near quiescence (McClintock et al., 1995; Froning et al., 2011; Gallo et al., 2019; Cherepashchuk et al., 2019). There is still no agreement on the nature of the additional component: it could be the jet, the accretion disk, the inner accretion flow, the gas stream, or some combination of the above. Different components may give similar contributions to the final spectrum, therefore additional information is needed to distinguish between them. Polarization of the optical radiation may provide such information.
@@
Indeed, optical and NIR radiation produced in various physical processes, including electron scattering of stellar radiation off the accretion disk/flow and synchrotron emission in the presence of the ordered magnetic field, can be polarized. The polarization degree (PD), polarization angle (PA), and their spectral properties are different for different processes, which makes polarimetry a powerful technique for studies of the physical mechanisms responsible for the optical and NIR emission production in BH X-ray binaries. A recent systematic study has shown, however, that in many quiescent BHs, the intrinsic optical polarization (corrected for interstellar contribution) is very small – in most cases (Kravtsov et al., 2022). On the other hand, several LMXBs show significant polarization during (or near) the quiescence: MAXI J1820+070 demonstrated high (up to ) PD with the blue spectrum and PA different from the jet direction, suggesting BH spin-orbit misalignment (Poutanen et al., 2022); A062000 showed 1% optical and NIR polarization in the quiescent state (Dubus et al., 2008; Russell et al., 2016).
@@
Residing in the X-ray quiescence, A062000 shows optical state changes: according to the Cantrell et al. (2008) classification, there are two different states of the optical activity –– passive and active. In the passive state, variations of the optical flux are consistent with the ellipsoidal variations, produced by the rotation of the tidally distorted optical companion. In the active state, the source is brighter and shows an aperiodic high-frequency variation usually called “flickering”.
@@
In this paper, we present the results of quasi-simultaneous high-precision optical polarimetric and multiwavelength (NIR to UV) photometric observations of A062000 during its passive quiescent state. The paper is organized as follows. In Sect. 2, we describe the observational data. In Sect. 3, we present the main results of our study: determination of the intrinsic optical polarization of A062000, its significant orbital variability, and rotation of the average intrinsic polarization angle with wavelength. In Sect. 4, we discuss possible physical mechanisms that can reproduce the observed behavior. Finally, in Sect. 5 we summarize our findings.
@@ 2 Data acquisition and analysis
2.1 Optical polarimetric observations
@@
High-precision optical polarimetric observations of A062000 were performed using broad-band BVR polarimeter DIPol-UF (Piirola et al., 2021), a visitor instrument installed at the 2.56 m Nordic Optical Telescope (NOT), Observatorio del Roque de los Muchachos, La Palma, Spain. Field stars used for the interstellar polarization estimation were observed with DIPol-2 (Piirola et al., 2014), mounted on the remotely-controlled 60 cm Tohoku telescope (T60) at Haleakala Observatory, Hawaii. Both polarimeters utilize a “double-image” design that effectively eliminates the polarization of the sky, even if it varies throughout the observations. The instrumental polarization is small () and well-calibrated by observing 20 unpolarized standard stars. The zero points of the PAs were determined by observing highly polarized standards HD 236928 and HD 25443. A more detailed description of the methods and calibrations can be found in Piirola et al. (2020) and Kravtsov et al. (2022), and references therein.
A062000 was observed during two nights between 2022 December 24–26 and 140 and 52 individual measurements of the Stokes parameters were made during the two nights, respectively. The first observation was continuous, 8 h long, and hence covered the whole orbital period of the binary, while the second observation covered only of the period. Hereafter we will refer to the average polarization measured during the first observation as orbit-averaged polarization. Each individual linear polarization measurement was obtained from four consecutive images with 50-s exposures taken at different half-wave plate positions, resulting in one polarization measurement per 3.3 min. To increase the signal-to-noise ratio, we split the data into bins so that each bin contains 10 individual measurements of the Stokes parameters. The errors of the normalized Stokes parameters and were computed as the standard errors of the weighted mean values within the bin. The Stokes parameters then were translated into the PD and PA
The uncertainty on the PD is equal to the uncertainty of the individual Stokes parameters, and uncertainty on the PA in radians was estimated as (Serkowski, 1962; Kosenkov et al., 2017).
2.2 Multiwavelength photometry
Quasi-simultaneous multiwavelength photometric observational campaign was organized on several telescopes (see Table 1). Near-infrared JHK photometry was made using EMIR wide-field imager (Garzón et al., 2022), installed on the 10.4 m Gran Telescopio Canarias (GTC), La Palma, Spain. Observations in the broadband SDSS-ugriz, Bessel V, and Bessel B filters were performed using IO:O instrument of the 2-m Liverpool Telescope (LT; Steele et al., 2004), La Palma, Spain. For all the instruments, basic data reductions such as bias and dark subtraction and flat fielding are done via the internal common pipelines.
2.3 Swift/UVOT
3 Results
3.1 Average intrinsic polarization
The observed optical polarization of a distant star is a combination of the intrinsic polarization of the object and interstellar (IS) polarization. The main source of the interstellar polarization (, ) is the optical dichroism of non-spherical interstellar dust particles aligned by Galactic magnetic field. To extract the intrinsic polarization (, ) of A062000 from the observed polarization (, ), we first estimated the IS polarization. We measured the polarization of the four close field stars (see Fig. 1 and Tab. 2) with similar parallaxes according to Gaia DR3 (Gaia Collaboration et al., 2021) and, taking into account the proximity of the star both in parallax and in angular distance on the image, we choose the polarization of the star Ref 1 as the best IS polarization estimate in the direction of A062000. Our estimate of the IS polarization of and (see Table 3) is consistent with the previous estimates in the band of and by Dubus et al. (2008) and and by Dolan (1976).
@@ 3.1 Average intrinsic polarization
Subtracting the interstellar component from the observed polarization, we find the average intrinsic polarization of A062000 to be –1.0% with changing from in the filter to in (see Table 3). Combining NIR polarimetric observations (Russell et al., 2016) corrected for IS polarization with our optical measurements, we see the continuous rotation of the intrinsic polarization vector with the wavelength from NIR to (see Fig. 3). The amplitude of the polarization vector rotation on the sky is (see Fig. 4). In contrast to the PA, PD barely changes with wavelength remaining at 1% level from to band.
@@
We note that although the determination of the intrinsic PA strongly depends on the accuracy of the IS polarization estimate, there are several additional reasons to believe that the intrinsic polarization estimate is close to its true value. First of all, our estimation of the IS polarization is consistent with the previous independent measurements (Dubus et al., 2008; Dolan, 1976). Furthermore, although the observed PD of A062000 in Dubus et al. (2008) -band observations was significantly higher than in this paper, subtracting our estimate of the IS polarization from their observed Stokes parameters, we get the PA of intrinsic polarization matching our value within a few degrees – this can be seen in Fig. 3, if one connects the origin to the corresponding observational points and compare the directions of the resulting vectors. This alignment of the intrinsic polarization vectors is unlikely to be coincidental and confirms our IS polarization estimate.
@@ 3.2 Orbital variability of the polarization
We folded the photometric and polarimetric observations of A062000 in -band (for which the S/N is the best) with the orbital period using the recent ephemeris of Cherepashchuk et al. (2019). We find statistically significant orbital variability of intrinsic polarization of the source in the filter (Fig. 5b,c). We plot the orbital profile of the observed polarization together with the optical light curve of A062000, obtained simultaneously (Fig. 5a). We see that the polarization degree is constant within the measurement errors during the first half of the orbit, and varies significantly in the second half. Variations of PD follow closely the variations of the R-band magnitudes. The absence of flickering in the photometric observations together with the visual magnitude of A062000 suggests that the source was in the passive quiescent state during our campaign (Cantrell et al., 2008).
@@ 3.3 Broadband spectrum
The observed NIR-to-UV spectral energy distribution (SED) of A062000 corrected for reddening assuming (Wu et al., 1983) is shown in Fig. 6. The key feature of the SED is the excess of UV photons, previously observed in the quiescent state (Froning et al., 2011; McClintock et al., 1995) – an extra source of UV radiation in addition to the optical companion is needed to reproduce the observed shape of the spectrum. The spectrum alone, however, does not allow unambiguous separation of the second component: even the simplest model with one additional blackbody is degenerate due to the uncertainty in the stellar spectrum normalization. Various different methods were used to estimate the contribution of non-stellar emission to the quiescent spectrum of A062000 (Marsh et al., 1994; Gelino et al., 2001; Harrison et al., 2007; Froning et al., 2011; Dinçer et al., 2018). Despite the fact that there is no complete consensus, it can be cautiously noted that most authors agree that the companion star dominates the entire NIR-to-optical range, with the contribution of additional component increasing towards the blue part of the spectrum. The additional component becomes comparable with the stellar flux only near band, where its contribution reaches 20–50%. Non-stellar emission spectrum can be described with the blackbody of temperature 9,000–11,000 K depending on the state of activity. The nature of the additional component is still under discussion, but most likely it corresponds to the brightest part of the accretion disk (either to the hot inner disk regions or to the bright spot formed at the impact point of the accretion stream). The bright spot is clearly present in the Doppler tomograms (Marsh et al., 1994; Shahbaz et al., 1994, 2004; Neilsen et al., 2008) and it is needed to explain the asymmetric light curves (Froning & Robinson, 2001; Cantrell et al., 2010; van Grunsven et al., 2017; Cherepashchuk et al., 2019). The contribution of the non-stellar radiation is variable on short (Haswell et al., 1993) and long (Cantrell et al., 2008) timescales. Our new observations are consistent with both interpretations (see Fig. 6 and Sect. 4).
@@ 4.1 The source of the polarization
The polarization measurements of optical light from the binary were obtained both in the quiescent state and in the outburst (Dolan, 1976, see the orange star in Fig. 3). The observed PD, PA, and their spectral properties were found to be consistent with the interstellar origin. Our IS polarization estimate is very close to the outburst polarization level of A062000, which confirms the above statement. Sub-per cent intrinsic optical polarization of A062000 is in line with that of the other X-ray binaries – most known sources observed during the outburst show low intrinsic polarization levels (XTE J1118+480, Schultz et al. 2004; MAXI J0637430, Kravtsov et al. 2019; MAXI J1820+070, Veledina et al. 2019; Kosenkov et al. 2020; SwiftJ 1727.81613, Kravtsov et al. 2023a). However, some BH LMXBs show a significant increase of intrinsic PD as they approach quiescence (e.g., MAXI J1820+070, Poutanen et al. 2022, and A062000, Dubus et al. 2008).
The orbit-averaged intrinsic polarization of A062000 in the quiescent state has nearly the same PD at all wavelengths from to band, yet PA shows a significant rotation over those wavelengths (Fig. 3). This makes A062000 the only source known to date for which such a strong PA dependence on the wavelength is observed. Such polarization behavior cannot be explained in terms of simple mechanisms of polarization production. Indeed, if the polarization is produced mostly by scattering, its PA should not depend on the wavelength.
We first study the case where Faraday rotation causes the observed dependence of PA on wavelength. We consider the scenario illustrated in Fig. 7. The unpolarized radiation of the companion star gets scattered by the accretion disk and the polarized scattered radiation has initial PA independent of the wavelength. After the scattering, the linearly polarized radiation propagates toward the observer through the magnetized plasma surrounding the accretion disk. The PA of the light propagating along the magnetic field lines experiences the Faraday rotation, resulting in the observed PA dependence on wavelength.
The PA of linearly polarized radiation propagating through the magnetized plasma rotates with the wavelength as:
where the rotation measure RM is defined through the integral along the line of sight
@@
We use Eq. (2) with parameters and RM to find the best-fit solution for the Faraday rotation of PA in our optical points and NIR points reported in Russell et al. (2016). We find the best-fit parameters and RM= rad m-2 (see Fig. 8). The latter value gives the relation between the line-of-sight optical depth and magnetic field (in Gauss) . For the realistic values of matter density in quiescence , we find G. The latter value is roughly consistent with the magnetic fields expected in the quiescent-state optically thin accretion flows (Wallace & Pe’er, 2021), which may be similar to those of optically thin plasma surrounding the disk (Fig. 7).
@@
We note, however, that the NIR data have been taken almost ten years prior to our optical observations, hence potential magnetic field variations may lead to the inconsistency of trends between the optical and NIR PA points. We indeed observe that our recent measurements (leftmost three points in Fig. 8) lie on a straight line with the slope differing from the general trend. If we apply Eq (2) only to our data, we obtain and RM= rad m-2 (see Fig. 8). This translates to , requiring almost seven times higher magnetic field strength for the Faraday rotation to occur in the optically thin plasma (). We also note that the Thomson scattering and Faraday effect cannot explain the complex dependence of PD on energy, with being higher than and . The sharp rise of polarization in a narrow wavelength range may be related to the atomic features.

### v2

# Optical and near-infrared polarization of the black hole X-ray binary A062000 in quiescence

@@
Optical and near-infrared polarization of the black hole X-ray binary A062000 in quiescence
We present simultaneous high-precision optical polarimetric and near-infrared (NIR) to ultraviolet (UV) photometric observations of the low-mass black hole X-ray binary A062000 in a quiescent state. Subtracting interstellar polarization, estimated from a sample of field stars, we derived the intrinsic polarization of A062000. We show that the intrinsic polarization degree (PD) varies with the orbital period with an amplitude of , at least in the band, where the signal-to-noise ratio of our observations is the best. This implies that some fraction of the optical polarization is produced by a scattering of stellar radiation off the matter that follows the black hole in its orbital motion. In addition, we see a rotation of the orbit-average intrinsic polarization angle (PA) with the wavelength from in the to in the band. All of the above, combined with the historical NIR-to-optical polarimetric observations, demonstrates the complex behavior of the average intrinsic polarization of A062000: the PA continuously rotates from the infrared to the blue band by in total, while the PD of remains nearly constant over the entire spectral range. The spectral dependence of the PA can be described by Faraday rotation with a rotation measure of rad m-2, implying a magnetic field of a few gauss in the plasma surrounding the black hole accretion disk. However, our preferred interpretation of the peculiar wavelength dependence is an interplay between two polarized components with different PAs. Polarimetric measurements in the UV range can help in distinguishing between these scenarios.
@@
A062000 is the prototypical black hole (BH) low-mass X-ray binary (LMXB). It was discovered during its 1975 outburst (Elvis et al., 1975), when its luminosity increased by more than a million times compared to the quiescent levels, reaching the Eddington luminosity. Since then, the object has been in a quiescent state, with its spectrum dominated by the emission of a K-type star orbiting a BH (McClintock & Remillard, 1986; van Grunsven et al., 2017) with an orbital period () of h. Nevertheless, additional sources of nonstellar origin (such as jet, accretion disk, inner hot accretion flow, and hot spot) were considered in the literature to explain the excess of radio, infrared (IR), and ultraviolet (UV) emission of A062000 near quiescence (McClintock et al., 1995; Muno & Mauerhan, 2006; Froning et al., 2011; Gallo et al., 2019; Cherepashchuk et al., 2019). Their contributions to the broadband spectrum are difficult to distinguish, especially in light of the alteration between accretion modes (passive, loop, and active modes; Cantrell et al. 2008). An additional source of information is needed to distinguish between different spectral models. Polarization of the optical radiation may provide such information.
@@
Optical and near-infrared (NIR) radiation produced in various physical processes, including the electron scattering of stellar radiation off the accretion disk or flow and synchrotron emission in the presence of an ordered magnetic field, can be polarized. The polarization degree (PD), polarization angle (PA), and their spectral properties are different for different processes, which makes polarimetry a powerful technique for studies of the physical mechanisms responsible for the optical and NIR emission production in BH X-ray binaries. A recent systematic study has shown, however, that in many quiescent BHs, the intrinsic optical polarization (corrected for interstellar contribution) is very small – in most cases (Kravtsov et al., 2022). On the other hand, several BH LMXBs show significant polarization during (or near) quiescence: MAXI J1820+070 had a high PD (up to ), with the blue spectrum and PA different from the jet direction, suggesting a BH spin-orbit misalignment (Poutanen et al., 2022); A062000 had 1% optical and NIR polarization in the quiescent state (Dubus et al., 2008; Russell et al., 2016).
@@
In X-ray quiescence, A062000 shows optical state changes: according to the Cantrell et al. (2008) classification, there are two different states of optical activity –– passive and active. In the passive state, variations in the optical flux are consistent with ellipsoidal variations produced by the rotation of the tidally distorted optical companion. In the active state, the source is brighter and shows an aperiodic high-frequency variation usually called “flickering.”
@@
In this paper we present the results of quasi-simultaneous high-precision optical polarimetric and multiwavelength (NIR to UV) photometric observations of A062000 during its passive quiescent state. The paper is organized as follows. In Sect. 2 we describe the observational data. In Sect. 3 we present the main results of our study: determination of the intrinsic optical polarization of A062000, its significant orbital variability, and the rotation of the average intrinsic PA with wavelength. In Sect. 4 we discuss possible physical mechanisms that can reproduce the observed behavior. Finally, in Sect. 5 we summarize our findings.
@@ 2 Data acquisition and analysis
2.1 Optical polarimetric observations
High-precision optical polarimetric observations of A062000 were performed using broadband BVR polarimeter DIPol-UF (Piirola et al., 2021), a visitor instrument installed at the 2.56 m Nordic Optical Telescope (NOT), Observatorio del Roque de los Muchachos, La Palma, Spain. Field stars used for the interstellar (IS) polarization estimation were observed with DIPol-2 (Piirola et al., 2014), mounted on the remotely controlled 60 cm Tohoku telescope (T60) at Haleakala Observatory, Hawaii. Both polarimeters utilize a “double-image” design that effectively eliminates the polarization of the sky, even if it varies throughout the observations. The instrumental polarization of both instruments is small () and was well calibrated by observing of 10–15 unpolarized standard stars. The zero points of the PAs were determined by observing highly polarized standards HD 236928 and HD 25443. A more detailed description of the methods and calibrations can be found in Piirola et al. (2020) and Kravtsov et al. (2022), and references therein.
A062000 was observed on two nights between 2022 December 24 and 26, and 140 and 52 individual measurements of the Stokes parameters were made on the two nights, respectively. The first observation was continuous, 8 h long, and hence covered the whole orbital period of the binary, while the second observation covered only 40% of the period. Hereafter we refer to the average polarization measured during the first observation as the orbit-averaged polarization. Each individual linear polarization measurement was obtained from four consecutive images with 50-s exposures taken at different half-wave plate positions, resulting in one polarization measurement per 3.3 min. To increase the signal-to-noise ratio, we split the data into bins such that each bin contains ten individual measurements of the Stokes parameters. The errors of the normalized Stokes parameters and were computed as the standard errors of the weighted mean values within the bin. The Stokes parameters then were translated into the PD () and PA (),
The uncertainty on the PD is equal to the uncertainty of the individual Stokes parameters, and uncertainty on the PA in radians was estimated as (Serkowski, 1962; Kosenkov et al., 2017). The phase-resolved PD was corrected for the bias caused by the low signal-to-noise ratio using the relation (Simmons & Stewart, 1985).
@@ 2.1 Optical polarimetric observations
2.2 Multiwavelength photometry
Quasi-simultaneous multiwavelength photometric observational campaign was organized on several telescopes (see Table 1). Near-infrared JHK photometry was made using EMIR wide-field imager (Garzón et al., 2022), installed on the 10.4 m Gran Telescopio Canarias (GTC), La Palma, Spain. Observations in the broadband SDSS-ugriz, Bessel V, and Bessel B filters were performed using IO:O instrument of the 2-m Liverpool Telescope (LT; Steele et al., 2004), La Palma, Spain. For all the instruments, basic data reductions such as bias and dark subtraction and flat fielding are done via the internal common pipelines. As the photometric standards, we used two stars with known Sloan Digital Sky Survey (SDSS) magnitudes. To obtain fluxes of the object from its magnitudes we used standard zero-points for SDSS and Johnson-Cousins systems (Fukugita et al., 1996; Bessell, 1979).
2.3 Swift/UVOT
3 Results
3.1 Average intrinsic polarization
@@
| A062000 | 1.33 | 0.11 | 0.11 | 1.52 | 0.07 | 0.07 | 1.33 | 0.04 | 0.04 | ||||||
The observed optical polarization of a distant star is a combination of the intrinsic polarization of the object and IS polarization. The main source of the IS polarization (, ) is the optical dichroism of nonspherical IS dust particles aligned by Galactic magnetic field. The observed polarization of A062000 is dominated by the IS component, as evidenced by the rough alignment (within 10°; see Fig. 1) of the observed PA of the source with those of the field stars. To extract the intrinsic polarization (, ) of A062000 from the observed polarization (, ), we first estimated the IS polarization. IS polarization usually depends on the distance to the source and may be different at different wavelengths. To study it, we observed four close field stars with similar Gaia DR3 (Gaia Collaboration et al., 2021) parallaxes in the BVR bands (see Table 2). The observations of the brightest field star (Ref 1) show very little wavelength dependence of IS polarization in the direction to A062000 in the optical band – it is consistent with being constant with wavelength at the level of better than (blue circle in Figs. 2 and 3). Thus, one zero point (, ) for IS polarization can be used for all three optical bands. Due to the possible dependence on distance, the exact position of this zero point on the -plane is uncertain. Different approaches can be used to set this zero point: one can use either the Stokes parameters of the closest in distance field star or the weighted average Stokes parameters of the group of field stars located at similar distances. We used the later approach to find (, ) based on the observations of the field stars Ref 1 – Ref 3, for which the accuracy of our observations was the best. The scatter in their Stokes parameters determines the systematic error on the IS polarization zero point (shown as red circle in Figs. 2 and 3). To find the intrinsic polarization of A062000, we subtracted the IS Stokes parameters (, ) from the observed Stokes parameters (, ) of the source. The statistical errors on the intrinsic Stokes parameters of A062000 were calculated as ; the term includes possible dependence of the IS polarization on wavelength. Then, Eq. (1) is used to calculate the PD and PA together with their statistical errors. The systematic error on the IS zero point translates to the systematic errors of the intrinsic PD and PA and .
@@
Subtracting the IS component from the observed polarization, we find the average intrinsic polarization of A062000 to be 0.8–1.0% with changing from in the filter to in (see Table 3). The fit with a constant to the intrinsic PA gives unacceptable fit with =10/2. Combining NIR polarimetric observations (Russell et al., 2016) corrected for IS polarization with our optical measurements, we see the continuous rotation of the intrinsic polarization vector with the wavelength from NIR to (see Fig. 3). The fit with a constant to all intrinsic PAs from NIR to B gives unacceptable fit with =32/6. The amplitude of the polarization vector rotation on the sky is (see Fig. 4). In contrast to the PA, PD barely changes with wavelength remaining at 1% level from to band (constant fit gives =5/6).
@@
We note that although the determination of the intrinsic PA strongly depends on the accuracy of the IS polarization estimate, there is an additional reason to believe that the intrinsic polarization estimate is close to its true value. Although the observed PD of A062000 in Dubus et al. (2008) -band observations was significantly higher than in this paper, subtracting our estimate of the IS polarization from their observed Stokes parameters, we get the PA of intrinsic polarization matching our value within a few degrees – this can be seen in Fig. 3, if one connects the origin to the corresponding observational points and compare the directions of the resulting vectors. This alignment of the intrinsic polarization vectors is unlikely to be coincidental and may appear naturally if the intrinsic PD of the source changes while the PA is stable.
@@ 3.2 Orbital variability of the polarization
We folded the photometric and polarimetric observations of A062000 in the band (for which the S/N is the best) with the orbital period using the recent ephemeris of Cherepashchuk et al. (2019). We find significant variability of intrinsic polarization of the source in the filter (Fig. 5b,c). We plotted the orbital profile of the observed polarization together with the optical light curve of A062000, obtained simultaneously (Fig. 5a). We see the pronounced peak in PD at the orbital phase of 0.75 and two minima at phases around 0.5 and 1.0. Similar peak in PD around phase 0.7 has been observed by Dolan & Tapia (1989). Such polarization behavior with minima in conjunctions and maxima in quadratures is expected and has been observed in other binary systems (discussed in detail in Sect. 4.3), and although our observations covered only 1.5 orbits, and the S/N for individual bins is at the threshold of significance, we can still cautiously state that the nature of observed polarization variability is more likely to be orbital rather than stochastic. The absence of flickering in the photometric observations together with the visual magnitude of A062000 suggests that the source was in the passive quiescent state during our campaign (Cantrell et al., 2008).
@@ 3.3 Broadband spectrum
The observed NIR-to-UV spectral energy distribution (SED) of A062000 in passive quiescent state, corrected for reddening assuming (Wu et al., 1983) is shown in Fig. 6. The key feature of the SED is the excess of UV photons, previously observed in the quiescent state (Froning et al., 2011; McClintock et al., 1995) – an extra source of UV radiation in addition to the optical companion is needed to reproduce the observed shape of the spectrum. The spectrum alone, however, does not allow an unambiguous separation of the second component: even the simplest model with one additional blackbody is degenerate due to the uncertainty in the stellar spectrum normalization. Various different methods were used to estimate the contribution of nonstellar emission to the quiescent spectrum of A062000 (Marsh et al., 1994; Gelino et al., 2001; Harrison et al., 2007; Froning et al., 2011; Dinçer et al., 2018). Despite the fact that there is no complete consensus, it can be cautiously noted that most authors agree that the companion star dominates the entire NIR-to-optical range in the passive quiescent state, with the contribution of additional component increasing toward the blue part of the spectrum. The additional component becomes comparable with the stellar flux only near the band, where its contribution reaches 20–50%. Nonstellar emission spectrum can be described with the blackbody of temperature 9000–11000 K depending on the state of activity. The nature of the additional component is still under discussion, but most likely it corresponds to the brightest part of the accretion disk (either the hot inner disk regions or the bright spot formed at the impact point of the accretion stream). The bright spot is clearly present in the Doppler tomograms (Marsh et al., 1994; Shahbaz et al., 1994, 2004; Neilsen et al., 2008) and it is needed to explain the asymmetric light curves (Froning & Robinson, 2001; Cantrell et al., 2010; van Grunsven et al., 2017; Cherepashchuk et al., 2019). The contribution of the nonstellar radiation is variable on short (Haswell et al., 1993) and long (Cantrell et al., 2008) timescales. Our new observations are consistent with both interpretations (see Fig. 6 and Sect. 4).
@@ 4.1 The source of the polarization
The first optical polarimetric observations of A062000 were performed during its 1975 outburst (Dolan, 1976, see the orange star in Fig. 3). The observed PD, PA, and their spectral properties were found to be consistent with the IS origin. Our IS polarization estimate is very close to the outburst polarization level of A062000, which confirms the above statement. Sub-percent intrinsic optical polarization of A062000 is in line with that of the other BH X-ray binaries – most known sources observed during the outburst show low intrinsic polarization levels (XTE J1118+480, Schultz et al. 2004; MAXI J0637430, Kravtsov et al. 2019; MAXI J1820+070, Veledina et al. 2019; Kosenkov et al. 2020; Swift J1727.81613, Kravtsov et al. 2023a; V404 Cyg, Tanaka et al. 2016; Kosenkov et al. 2017; GX 3394, Mastroserio et al. 2025; LMC X-3, Boyd et al. 2001). However, the intrinsic PD of some BH LMXBs increases significantly as they approach quiescence (e.g., MAXI J1820+070, Poutanen et al. 2022, and A062000, Dubus et al. 2008). Intrinsic optical polarization is also detected in some NS X-ray binaries (Aql X-1, Charles et al. 1980; Cyg X-2, Koch Miramond & Naylor 1995; Sco X-1, Schultz et al. 2004; 4U 0614+091, Baglio et al. 2014; XTE J1709–267, Higgins et al. 2019).
The orbit-averaged intrinsic polarization of A062000 in the quiescent state has nearly the same PD at all wavelengths from the to band, yet PA shows a significant rotation over those wavelengths (Fig. 3). This makes A062000 the only source known to date for which such a strong PA dependence on the wavelength is observed. Such polarization behavior cannot be explained in terms of simple mechanisms of polarization production. Indeed, if the polarization is produced mostly by scattering, its PA should not depend on the wavelength.
We first studied the case in which the observed dependence of PA on wavelength is caused by Faraday rotation. We considered the scenario illustrated in Fig. 7. The unpolarized radiation of the companion star gets scattered by the accretion disk and the polarized scattered radiation has initial PA independent of the wavelength. After the scattering, the linearly polarized radiation propagates toward the observer through the magnetized plasma surrounding the accretion disk. The PA of the light propagating along the magnetic field lines experiences the Faraday rotation, resulting in the observed PA dependence on wavelength.
The PA of linearly polarized radiation propagating through the magnetized plasma rotates with the wavelength as
where the rotation measure (RM) is defined through the integral along the line of sight:
@@
We used Eq. (2) with parameters and RM to find the best-fit solution for the Faraday rotation of PA in our optical points and NIR points reported in Russell et al. (2016). We find the best-fit parameters and RM= rad m-2 (see Fig. 8). The latter value gives the relation between the line-of-sight optical depth and magnetic field (in gauss) . For the realistic values of matter density in quiescence , we find G. This value is roughly consistent with the magnetic fields expected in the quiescent-state optically thin accretion flows (Wallace & Pe’er, 2021), which may be similar to those of optically thin plasma surrounding the disk (Fig. 7).
@@
We note, however, that the NIR data were taken almost ten years prior to our optical observations, and hence potential magnetic field variations may lead to the inconsistency of trends between the optical and NIR PA points. We indeed observe that our recent measurements (leftmost three points in Fig. 8) lie on a straight line with the slope differing from the general trend. If we apply Eq (2) only to our data, we obtain and RM= rad m-2 (see Fig. 8). This translates to , requiring almost seven times higher magnetic field strength for the Faraday rotation to occur in the optically thin plasma ().